OpenStudy (anonymous):

If f(x)=3x^2-8 and g(x) = 2x, evaluate f(x)/g(x) for x=2 A. 3 B.13/3 C.1 D.1/3

OpenStudy (anonymous):

please answer fast i got 2 mins

OpenStudy (anonymous):

c

OpenStudy (anonymous):

f(2)=(3*4-8)=4 g(2)=2*2=4 f(2)/g(2)=4/4=1

OpenStudy (anonymous):

@deidara3 ,ok?

OpenStudy (anonymous):

alright thx!

OpenStudy (anonymous):

sorry is 3*x^2 or (3x)^2 ?

OpenStudy (campbell_st):

so are you doing a quiz... and asking for answers..?

OpenStudy (anonymous):

@campbell_st No, It's a homework question

OpenStudy (campbell_st):

so its online... is that why there's a time limit...

OpenStudy (anonymous):

what are u talkin about..?

OpenStudy (campbell_st):

your initial comment... pls answer fast.. I have 2 mins

OpenStudy (anonymous):

nothing,i think it is 3*x^2

OpenStudy (campbell_st):

makes me wonder how learning occurs... if you just ask people for homework answers you can't do...

OpenStudy (anonymous):

because i got to go to sleep and i don't have time gotta be done tomorrow

OpenStudy (campbell_st):

so whats the penalty for not completing a question on homework..?

OpenStudy (anonymous):

suspension... are u an admin or what?

OpenStudy (campbell_st):

nope... just curious... i'm in 8th grade... if I can't do a question on homework... I leave it and then see the teacher before class to explain it

OpenStudy (anonymous):

Ok...

OpenStudy (campbell_st):

but you just seem to ask for answers... which for me is cheating sorry to say that

OpenStudy (anonymous):

is alright but whats the worst thing that could happen?

OpenStudy (campbell_st):

how can you be penalised if you don't understand the concept...

OpenStudy (anonymous):

C.1